OpenStudy (anonymous):

help please Addison and Caden prefer tetra fish to goldfish. Addison buys 5 neon tetras and 4 cardinal tetras for $15.50. Caden buys 6 neon tetras and 10 cardinal tetras for $26.40. A. Write a system of equations that can be used to find n, the cost of 1 neon tetra, and c, the cost of 1 cardinal tetra. B.What is the cost of 1 neon tetra?

OpenStudy (owlcoffee):

Here, you have to set up a system of equalities that allows us to model the situation. Let's first model what Addison bought; this being 5 neon tetras (n) and 4 cardinal tetras (c) and paid 15.5$ so therefore, a model would be: \[5n +4c=15,50\] That will be the first equation, now for Cadens model; she bought 6 neon tetras and 10 cardinal tetra, and paid 26.40$ which means that the model for her purchse is: \[6n +10c=26,40\] Now, in order to find the prices for each type of fish we will then build the system of equation, and it will be a 2x2 type, since we have two variables and two equations: \[5n+4c=15,50\] \[6n+10c=26,40\]

OpenStudy (owlcoffee):

Now you can solve that using any method you have been taught.

OpenStudy (anonymous):

thank youuuu

OpenStudy (anonymous):

yay @Tommynaiter

OpenStudy (anonymous):

You need to setup 2 equations with 2 variables. You get the following: \[5n+4c=15.50\]\[6n+10c=26.40\] Here is n neon and c is cardinal. Do you see where the equations come from?

OpenStudy (anonymous):

yes but i have trouble finding the price

OpenStudy (anonymous):

Alright. So do you know the method substitution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So you want to use that.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

So lets do it in small steps. Start by isolating n in the first equation. \[5n+4c=15.50\]\[n=?\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you know how to isolate n?

OpenStudy (anonymous):

you divide

OpenStudy (anonymous):

First we subtract 4c. \[5n+4c=15.50\]\[5n=15.50-4c\] Now we divide with 5 \[n=3.1-\frac{4}{5}c\]

OpenStudy (anonymous):

subtract 4c

OpenStudy (anonymous):

Yeap, it looks like you know how to solve n.

OpenStudy (anonymous):

So the next step is to insert the value for n into the 2nd equation.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We had the 2nd equation: \[6n+10c=26.40\] Insert the values for n: \[6*(3.1-\frac{4}{5}c)+10c=26.40\]

OpenStudy (anonymous):

Do you follow what I just did?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good, so now we want to isolate c. Then we have the price for c.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

you divide

OpenStudy (anonymous):

right

OpenStudy (anonymous):

There is a couple of steps. First I would get rid of the parentheses.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[6∗(3.1−\frac{4}{5}c)+10c=26.40\]\[6*3.1=18.6\]\[6*(-\frac{4}{5}c)=-\frac{24}{5}c\] So you get \[18.6-\frac{24}{5}c+10c=26.40\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So we continue to solve c. Subtract 18.6 \[-\frac{24}{5}c+10c=7.8\] add the c's together \[\frac{26}{5}c=7.8\] multiply with 5 \[26c=39\] divide with 26 \[c=1.5\]

OpenStudy (anonymous):

Do you understand how I solved c?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good, so now that we know the price for cardinal. We just insert the price into either one of the equations and find a value for n. \[5n+4c=15.50\] insert \[c=1.5\] \[5n+4*1.5=15.50\]\[5n+6=15.50\] subtract 6 \[5n=9.50\] divide with 5 \[n=1.90\]

OpenStudy (anonymous):

So the price for the cardinal (c) is 1.5 and the price for neon (n) is 1.9

OpenStudy (anonymous):

ok thank youuuu

OpenStudy (anonymous):

Yea, I hope you understand it :) else just ask

OpenStudy (anonymous):

I am always willing to help, if you'd like you can follow me to see if im online :)

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

You're welcome, have a good day @ameliagrace417 :D

OpenStudy (anonymous):

you tooo @Tommynaiter :)